Each edge of a cube is a-2 units long . Which shows its volume in cubic units
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

It's (a-2)*(a-2)*(a-2) cubic units

Answer by stanbon(75887)

Each edge of a cube is a-2 units long . Which shows its volume in cubic units
-----
Volume = (a-2)^3 = a^3+3a^2(-2)+3a(-2)^2+(-2)^3
-----
= a^3 -6a^2+12a - 8
=========================
Cheers,
Stan H.
===============